OpenStudy (anonymous):

d. passes through (1, 2) with slope -5/3

OpenStudy (anonymous):

u want its equation? y= -5/3x+11/3

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Hi there! let's make a line! First, do you know the point-slope form of a line?

OpenStudy (anonymous):

y-2=-5/3(x-2)